# More is Less: The Pitfalls of Multi-Model Synthetic Preference Data in DPO Safety Alignment

### v1

@@ Abstract
Keywords Alignment  $\cdot$ Synthetic Data  $\cdot$ Safety  $\cdot$ Large Language Models
@@ 1 Introduction
Large Language Models (LLMs) have emerged as general artificial intelligence models that tackle most language tasks in zero or a few shots. To ensure that these models behave as expected, they must be aligned with human preferences. Alignment involves guiding language models to adhere to human values and prevent harmful outputs, while enhancing capabilities in areas such as question answering, coding, math and reasoning. Reinforcement Learning from Human Feedback (RLHF) is a widely used alignment method that uses human preference data as a reward signal to guide model behavior. However, due to the computational expense, complexity, and training instability associated with traditional RL methods such as Proximal Policy Optimization (PPO)[1], researchers have developed Direct Preference Optimization (DPO) [2], a simpler method that achieves comparable alignment results. Building upon DPO’s success, several variants have emerged, including IPO [3], CPO [4], ORPO [5], KTO [6], SimPO [5] and others, each with different training objectives.
A fundamental challenge in alignment lies in creating high-quality preference datasets. This process involves data collection and preference labeling. For data collection, preference datasets primarily utilize synthetic data generated by a single model or multiple models. Examples of single-model generated datasets include HH-RLHF/harmless-base [7] (using a 52B context-distilled language model) and BeaverTails [8] (using a 7B Alpaca model). Multi-model datasets include UltraFeedback [9] (sampling from 17 different models), SafeRLHF [10] (utilizing 7B, 8B, and 70B Alpaca models), and Chatbot Arena [11] (sourcing from 20 distinct models). The second key element of alignment is preference labeling. Beyond human annotation, researchers have explored AI-assisted labeling methods such as RLAIF and Constitutional AI [12], which use LLM feedback as reward signals. Furthermore, there are rule-based methods such as rule-based reward models [13] or simple heuristics that designate stronger models’ responses or human-chosen responses as chosen examples and self-generated responses as rejected examples [14, 15].
@@ 2 Related Work
$r(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi_{\rho}(\mathbf{y}\mid\mathbf{x})}{\pi_{\mathrm{rel}}(\mathbf{y}\mid\mathbf{x})}$
@@
While these alignment methods aim to create safe and helpful models, researchers have demonstrated various ways to circumvent these safety measures through “jailbreaking" attacks. These attacks can be categorized into three major groups: manually designed, model-generated, and adversarial optimization-based attacks. Manually designed attacks exploit weaknesses in LLMs using techniques like prefix injection [16], base64 encoding [17], and in-context learning attacks [18]. Model-generated attacks leverage LLMs themselves to create effective jailbreak prompts, using methods like PAIR [19] and PMA [20]. Adversarial optimization-based attacks employ sophisticated techniques like GCG [21] and HGA [22] to develop transferable and effective attack prompts. We use the advbench dataset from [21] to evaluate jailbreaking attack success rate (ASR) as our primary safety metric.
@@
Recent approaches to synthetic preference data creation can be categorized into single-model and multi-model generation methods. Single-model generation uses only one language model as the sole generator of candidate outputs. By adjusting high-entropy sampling parameters, a single model can generate a diverse set of answers to the same prompt. Human annotators (or an automated judge) then compare these outputs to create preference labels. For example, the well-known HH-RLHF dataset [7] primarily uses a 52B context-distilled language model to generate multiple candidate responses per prompt for further training. Similarly, the BeaverTails safety alignment dataset [8] was built by prompting a 7B model (Alpaca) to produce multiple answers per query. In contrast, multi-model generation uses different models as sources of candidate responses for each prompt, increasing diversity through varied architectures or training backgrounds. UltraFeedback [9] samples four out of 17 different models per prompt to enhance diversity. SafeRLHF [10] uses multiple Alpaca-derived models of varying sizes (7B, 8B, and 70B) to generate responses, focusing on prompt diversity through contextual augmentation. Chatbot Arena [11] leverages this approach by sourcing responses from a pool of 20 distinct models, including strong online models like GPT-4 and Claude-V1 as well as many open-source models. Another common approach is to use the responses from a more powerful model directly as the chosen answer, while using target model’s or weaker model’s response as rejected answers. For example, Intel’s orca_dpo_pairs dataset employs GPT-4 generated responses as the chosen answers and uses responses from the LLaMA-13B model as rejected answers.
@@
Reward hacking [23] occurs when an agent exploits flaws in a reward function to maximize rewards without achieving the intended objectives. In the context of LLM alignment [24], the designer of a model often has some true reward function $R_{t}$ they want the model to optimize for, and some proxy reward function $R_{p}$ they actually optimize for. Depending on the optimization method, $R_{p}$ can be either explicit (as in PPO), or implicit (as in DPO). One example of data induced reward hacking is that models may learn that longer responses [25] or specific formatting patterns (e.g., bullet points, emojis [26], examples [25]) correlate with higher human preferences, leading them to optimize for these irrelevant features. Likewise, frequent use of safety-related keywords (e.g., “I apologize," “I cannot") can become spurious markers for higher preference [27]. Recent approaches like RRM [28] attempt to counteract these effects by augmenting and rebalancing preference data to disrupt spurious features; however, achieving robust alignment without inadvertently rewarding superficial traits remains an ongoing challenge.
@@ 3.1 Experimental Setup
In our experiments, we test 6 prevalent models from 3 different model families: LLaMA3.1-8B, LLaMA2-7B, Mistral-7B-v0.3, Mistral-7B-v0.1, Qwen2.5-7B, and Qwen2-7B. Prior to preference training, we conduct supervised fine-tuning (SFT) using a combined dataset of 50K samples from HuggingFaceH4/ultrafeedback_binarized [9] for general capabilities and 50K PKU-Alignment/PKU-SafeRLHF [10] samples for safety fine-tuning. SFT training details can be found in Appendix B.1.
@@
For our main experiments, we use one general preference dataset, HuggingFaceH4/ultrafeedback_binarized [9], and one safety-related dataset, PKU-Alignment/PKU-SafeRLHF [10]. Additional safety experiments with Anthropic/hh-rlhf/harmless-base [7] are presented in the Appendix A.3. We construct preference data using a uniform set of 10,000 prompts drawn from these datasets, exploring different approaches to create preference data pairs (chosen vs. rejected responses). We investigate common single-model and multi-model data generation approaches:
@@
    • Self+RM: The target model generates multiple responses, and a reward model picks the highest-scoring as chosen and lowest-scoring as rejected.
@@
We evaluate general performance on five widely used tasks: ARC [29], HellaSwag [30], Winogrande [31] for commonsense reasoning, MMLU [32] for multi-task language understanding, and TruthfulQA [33] for human falsehood mimic.
For safety evaluation, we use AdvBench introduced in [21] as the jailbreaking test set and measure attack success rate (ASR) as the safety metric. GPT-4o is used as a judge to evaluate the harmfulness of model responses, scoring them on a scale from 1 (the least harmful) to 5 (the most harmful). Detailed safety evaluation instruction for GPT-4o is presented in Appendix B.4. Responses rated 5 are classified as jailbroken. The ASR is calculated as the percentage of responses with a score of 5 among 100 samples.
@@ 3.3 Multi-Model Generated Data Fails in Safety Alignment
We also investigate whether multi-model response pool could enhance safety by leveraging diverse responses from different models. We collect responses from 6 peer models and all 11 available models, forming the “peers pool" and “all models pool", respectively. A reward model then selects the best and worst responses from the pool to serve as chosen and rejected responses (see Appendix B.3 for details). As shown in Figure 1, both Peer+RM and All+RM perform better than direct pairing methods. However, they still underperform compared to Self+RM alone.
The performance hierarchy that emerges from our experiments is clear and consistent: Self+RM outperforms multi-model response pool approaches (All+RM / Peer+RM), which in turn perform better than direct pairing methods (GPT4o+Self / Stronger+Self / HC+Self). This ordering suggests that the effectiveness of safety alignment is strongly influenced by the distributional similarity between preference data and the model’s own output distribution. While mixing data sources might intuitively seem beneficial for diversity, in practice it appears to dilute the effectiveness of safety training by forcing the model to learn from examples that do not match its own capabilities and behavioral patterns.
@@ 3.4 Analysis: Why Multi-Model Data Fails in Safety Alignment
We investigate why some preference data creation methods fail at safety alignment by comparing two representative approaches — one from multi-model generation and one from single-model generation: (1) GPT4o+Self, which pairs seemingly “better responses" with self-generated responses; and (2) Self+RM, which solely relies on self-generated data ranked by a reward model. Through this comparison, we aim to understand why methods using external “better responses" often lead to poor safety performance compared to self-alignment approaches.
@@
Our results reveal a non-linear relationship between linear separability and safety performance. When classification loss is extremely low (high separability), models appear to learn superficial patterns that easily distinguish between chosen and rejected responses without capturing meaningful safety concepts. This suggests that when the distinction is too obvious or based on simple stylistic features, models fail to learn substantive safety preferences and instead exploit these superficial patterns. Conversely, when classification loss is high (low separability), chosen and rejected responses become difficult to distinguish. While this avoids learning superficial patterns, it indicates that the safety preference signal may be too weak or inconsistent – the chosen and rejected responses might have minimal safety-relevant differences or even contradictory labels. This explains why such cases still don’t achieve optimal safety results despite high-separability. Our data reveals that optimal safety performance (lowest ASR) occurs at moderate linear separability levels, suggesting a “sweet spot" where the distinction between chosen and rejected responses is clear enough to provide a learning signal but not so obvious that it can be captured by simple stylistic features. This moderate separability appears to force models to learn more nuanced safety-relevant features, leading to better generalization and more robust safety alignment.
@@ 4 Discussion
Existing results have shown that “Length Overfitting" in DPO Training causes models to generate excessively long outputs by overemphasizing the “longer = better" heuristic [25, 5]. Hence, we conducted an additional analysis on the length difference between the chosen and rejected responses. As shown in Figure 3, we found that the average length of self-generated responses based on SafeRLHF prompts is significantly longer than original responses, while the response length in general dataset Ultrafeedback Binarized is lower. Interestingly, only the safety-related task shows a significant performance drop, while performance on general tasks remains comparable.
@@ A.2 Full Table of Safety Evaluation
Table 4 presents all safety evaluation results across multiple metrics: GPT-ASR (the primary metric discussed in the paper), Keyword ASR (KW-ASR) as described in [21], and GPT-Score (on a scale of 1 to 5). Our evaluation encompasses various data pairing approaches, all using an identical set of 10,000 prompts from the SafeRLHF dataset.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have emerged as general artificial intelligence models that tackle most language tasks in zero or a few shots. To ensure that these models behave as expected, they must be aligned with human preferences. Alignment involves guiding language models to adhere to human values and prevent harmful outputs, while enhancing capabilities in areas such as question answering, coding, math and reasoning. Reinforcement Learning from Human Feedback (RLHF) is a widely used alignment method that uses human preference data as a reward signal to guide model behavior. However, due to the computational expense, complexity, and training instability associated with traditional RL methods such as Proximal Policy Optimization (PPO)(Schulman et al., 2017), researchers have developed Direct Preference Optimization (DPO) (Rafailov et al., 2024), a simpler method that achieves comparable alignment results. Building upon DPO’s success, several variants have emerged, including IPO (Azar et al., 2023), CPO (Xu et al., 2024), ORPO (Hong et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024) and others, each with different training objectives.
A fundamental challenge in alignment lies in creating high-quality preference datasets. This process involves data collection and preference labeling. For data collection, preference datasets primarily utilize synthetic data generated by a single model or multiple models. Examples of single-model generated datasets include HH-RLHF/harmless-base (Bai et al., 2022a) (using a 52B context-distilled language model) and BeaverTails (Ji et al., 2023) (using a 7B Alpaca model). Multi-model datasets include UltraFeedback (Cui et al., 2024) (sampling from 17 different models), SafeRLHF (Dai et al., 2023) (utilizing 7B, 8B, and 70B Alpaca models), and Chatbot Arena (Chiang et al., 2024) (sourcing from 20 distinct models). The second key element of alignment is preference labeling. Beyond human annotation, researchers have explored AI-assisted labeling methods such as RLAIF and Constitutional AI (Bai et al., 2022b), which use LLM feedback as reward signals. Furthermore, there are rule-based methods such as rule-based reward models (Mu et al., 2024) or simple heuristics that designate stronger models’ responses or human-chosen responses as chosen examples and self-generated responses as rejected examples (Rosset et al., 2024; Burns et al., 2023).
@@ 2 Related Work
$r(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi_{\theta}(\mathbf{y}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})}$
@@
While these alignment methods aim to create safe and helpful models, researchers have demonstrated various ways to circumvent these safety measures through “jailbreaking” attacks. These attacks can be categorized into three major groups: manually designed, model-generated, and adversarial optimization-based attacks. Manually designed attacks exploit weaknesses in LLMs using techniques like prefix injection (Wei et al., 2024), base64 encoding (Shayegani et al., 2023), and in-context learning attacks (Li et al., 2023). Model-generated attacks leverage LLMs themselves to create effective jailbreak prompts, using methods like PAIR (Chao et al., 2023) and PMA (Shah et al., 2024). Adversarial optimization-based attacks employ sophisticated techniques like GCG (Zou et al., 2023) and HGA (Liu et al., 2023) to develop transferable and effective attack prompts. We use the advbench dataset from (Zou et al., 2023) to evaluate jailbreaking attack success rate (ASR) as our primary safety metric.
@@
Recent approaches to synthetic preference data creation can be categorized into single-model and multi-model generation methods. Single-model generation uses only one language model as the sole generator of candidate outputs. By adjusting high-entropy sampling parameters, a single model can generate a diverse set of answers to the same prompt. Human annotators (or an automated judge) then compare these outputs to create preference labels. For example, the well-known HH-RLHF dataset (Bai et al., 2022a) primarily uses a 52B context-distilled language model to generate multiple candidate responses per prompt for further training. Similarly, the BeaverTails safety alignment dataset (Ji et al., 2023) was built by prompting a 7B model (Alpaca) to produce multiple answers per query. In contrast, multi-model generation uses different models as sources of candidate responses for each prompt, increasing diversity through varied architectures or training backgrounds. UltraFeedback (Cui et al., 2024) samples four out of 17 different models per prompt to enhance diversity. SafeRLHF (Dai et al., 2023) uses multiple Alpaca-derived models of varying sizes (7B, 8B, and 70B) to generate responses, focusing on prompt diversity through contextual augmentation. Chatbot Arena (Chiang et al., 2024) leverages this approach by sourcing responses from a pool of 20 distinct models, including strong online models like GPT-4 and Claude-V1 as well as many open-source models. Another common approach is to use the responses from a more powerful model directly as the chosen answer, while using target model’s or weaker model’s response as rejected answers. For example, Intel’s orca_dpo_pairs dataset employs GPT-4 generated responses as the chosen answers and uses responses from the LLaMA-13B model as rejected answers.
@@
Reward hacking (Skalse et al., 2022) occurs when an agent exploits flaws in a reward function to maximize rewards without achieving the intended objectives. In the context of LLM alignment (Weng, 2024), the designer of a model often has some true reward function $R_{t}$ they want the model to optimize for, and some proxy reward function $R_{p}$ they actually optimize for. Depending on the optimization method, $R_{p}$ can be either explicit (as in PPO), or implicit (as in DPO). One example of data induced reward hacking is that models may learn that longer responses (Park et al., 2024) or specific formatting patterns (e.g., bullet points, emojis (Zhang et al., 2024), examples (Park et al., 2024)) correlate with higher human preferences, leading them to optimize for these irrelevant features. Likewise, frequent use of safety-related keywords (e.g., “I apologize,” “I cannot”) can become spurious markers for higher preference (Qi et al., 2024). Recent approaches like RRM (Liu et al., 2024) attempt to counteract these effects by augmenting and rebalancing preference data to disrupt spurious features; however, achieving robust alignment without inadvertently rewarding superficial traits remains an ongoing challenge.
@@ 3.1 Experimental Setup
In our main experiments, we test 6 prevalent models from 3 different model families: LLaMA3.1-8B, LLaMA2-7B, Mistral-7B-v0.3, Mistral-7B-v0.1, Qwen2.5-7B, and Qwen2-7B. Prior to preference training, we conduct supervised fine-tuning (SFT) using a combined dataset of 50K samples from HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2024) for general capabilities and 50K PKU-Alignment/PKU-SafeRLHF (Dai et al., 2023) samples for safety fine-tuning. SFT training details can be found in Appendix B.1.
@@
For our main experiments, we use one general preference dataset, HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2024), and one safety-related dataset, PKU-Alignment/PKU-SafeRLHF (Dai et al., 2023). We construct preference data using a uniform set of 10,000 prompts drawn from these datasets, exploring different approaches to create preference data pairs (chosen vs. rejected responses). We investigate common single-model and multi-model data generation approaches:
@@
    • Self+RM: The target model generates multiple responses, and a reward model [^1] In all our experiments, we use the same reward model: OpenAssistant/reward-model-deberta-v3-large-v2 picks the highest-scoring as chosen and lowest-scoring as rejected.
@@
We evaluate general performance on five widely used tasks: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2019) for commonsense reasoning, MMLU (Hendrycks et al., 2021) for multi-task language understanding, and TruthfulQA (Lin et al., 2022) for human falsehood mimic.
For safety evaluation, we use AdvBench introduced in (Zou et al., 2023) as the jailbreaking test set and measure attack success rate (ASR) as the safety metric. GPT-4o is used as a judge to evaluate the harmfulness of model responses, scoring them on a scale from 1 (the least harmful) to 5 (the most harmful). Detailed safety evaluation instruction for GPT-4o is presented in Appendix B.5. Responses rated 5 are classified as jailbroken. The ASR is calculated as the percentage of responses with a score of 5 among 100 samples.
@@ 3.3 Multi-Model Generated Data Fails in Safety Alignment
We also investigate whether multi-model response pool could enhance safety by leveraging diverse responses from different models. We collect responses from 6 peer models and all 11 available models, forming the “peers pool” and “all models pool”, respectively. A reward model then selects the best and worst responses from the pool to serve as chosen and rejected responses (see Appendix B.4 for details). As shown in Figure 1, both Peer+RM and All+RM perform better than direct pairing methods. However, they still underperform compared to Self+RM alone.
The performance hierarchy that emerges from our experiments is clear and consistent: Self+RM outperforms models’ responses pool approaches (All+RM / Peer+RM), which in turn perform better than direct pairing methods (GPT4o+Self / Stronger+Self / HC+Self). This ordering suggests that the effectiveness of safety alignment is strongly influenced by the distributional similarity between preference data and the model’s own output distribution. While mixing data sources might intuitively seem beneficial for diversity, in practice it appears to dilute the effectiveness of safety training by forcing the model to learn from examples that do not match its own capabilities and behavioral patterns.
@@ 3.4 Analysis: Why Multi-Model Data Fails in Safety Alignment
We investigate why some preference data creation methods fail at safety alignment by comparing two representative approaches — one from multi-model generation and one from single-model generation: (1) GPT4o+Self, which pairs seemingly “better responses” with self-generated responses; and (2) Self+RM, which solely relies on self-generated data ranked by a reward model. Through this comparison, we aim to understand why methods using external “better responses” often lead to poor safety performance compared to self-alignment approaches.
@@
Our results reveal a non-linear relationship between linear separability and safety performance. When classification loss is extremely low (high separability), models appear to learn superficial patterns that easily distinguish between chosen and rejected responses without capturing meaningful safety concepts. This suggests that when the distinction is too obvious or based on simple stylistic features, models fail to learn substantive safety preferences and instead exploit these superficial patterns. Conversely, when classification loss is high (low separability), chosen and rejected responses become difficult to distinguish. While this avoids learning superficial patterns, it indicates that the safety preference signal may be too weak or inconsistent – the chosen and rejected responses might have minimal safety-relevant differences or even contradictory labels. This explains why such cases still don’t achieve optimal safety results despite high-separability. Our data reveals that optimal safety performance (lowest ASR) occurs at moderate linear separability levels, suggesting a “sweet spot” where the distinction between chosen and rejected responses is clear enough to provide a learning signal but not so obvious that it can be captured by simple stylistic features. This moderate separability appears to force models to learn more nuanced safety-relevant features, leading to better generalization and more robust safety alignment.
3.5 Generalizability of Findings
We further validate the generalizability of our core finding across four key dimensions: datasets, model scales, optimization methods, and safety evaluation metrics. Different dataset validation: Additional experiments on the HH-RLHF harmful dataset further confirm that the benefits of Self+RM are not dataset-specific (Appendix A.3). Scale robustness: Experiments on larger models (Llama-2-13B and Qwen2.5-14B) confirm that Self+RM maintains superior safety performance regardless of model capacity, consistently achieving lower ASRs compared to multi-model approaches (detailed results in Appendix A.4). Method generalizability: Our conclusions extend beyond DPO to other direct preference optimization methods like Implicit Preference Optimization (IPO) (Azar et al., 2023) (Appendix A.5) and online DPO settings, where Self+RM continues to outperform multi-model strategies across multiple training iterations(Appendix A.6). Broader safety evaluation: The superiority of self-generated preference data extends beyond jailbreak resistance to encompass toxicity mitigation and fairness, as evidenced by Self+RM achieving the lowest toxicity scores on the TOXIGEN dataset (Appendix A.7). These results collectively demonstrate the broad applicability of our findings.
@@ 4 Discussion
Existing results have shown that “Length Overfitting” in DPO Training causes models to generate excessively long outputs by overemphasizing the “longer = better” heuristic (Park et al., 2024; Meng et al., 2024). Hence, we conducted an additional analysis on the length difference between the chosen and rejected responses. As shown in Figure 3, we found that the average length of self-generated responses based on SafeRLHF prompts is significantly longer than original responses, while the response length in general dataset Ultrafeedback Binarized is lower. Interestingly, only the safety-related task shows a significant performance drop, while performance on general tasks remains comparable.
@@ A.2 Full Table of Safety Evaluation
Table 4 presents all safety evaluation results across multiple metrics: GPT-ASR (the primary metric discussed in the paper), Keyword ASR (KW-ASR) as described in (Zou et al., 2023), and GPT-Score (on a scale of 1 to 5). Our evaluation encompasses various data pairing approaches, all using an identical set of 10,000 prompts from the SafeRLHF dataset.
@@ A.3 Additional Experiments on HH-RLHF Dataset
A.4 Additional Experiments on Larger Models
To investigate the generalizability of our findings to larger models, we conducted experiments on two larger model variants: Llama-2-13B and Qwen2.5-14B.
Table 5 presents the Attack Success Rate (ASR) results for both larger models across different preference data creation strategies. The evaluation uses the same set of 10,000 prompts from the SafeRLHF dataset to ensure consistency with our main experiments.
Model Method ASR
Llama-2-13B HC+Self 0.46
Stronger+Self 0.32
GPT4o+Self 0.23
Self+RM 0.13
Qwen2.5-14B HC+Self 0.47
Stronger+Self 0.55
GPT4o+Self 0.26
Self+RM 0.14
Table 5: Attack Success Rate comparison for larger models across different preference data creation methods.
As shown in Table 5, both larger models demonstrate the same relative trend across data sourcing methods: Self+RM consistently yields the lowest ASR, while HC+Self and GPT-4o/Stronger+Self produce higher ASR values. These results demonstrate that our core findings on the safety impact of data curation hold across model scales, further reinforcing our conclusion.
A.5 Additional Experiments on Other Direct Preference Optimization Methods
To validate the generalizability of our findings beyond DPO, we conducted experiments using other direct preference optimization methods. Specifically, we evaluated Implicit Preference Optimization (IPO) on Llama-3.1-8B, employing the same data generation strategies as in our DPO experiments to ensure fair comparison.
Table 6 presents the Attack Success Rate (ASR) results for IPO across different preference data creation methods.
Method ASR
HC+Self 0.47
GPT4o+Self 0.19
Stronger+Self 0.09
Human-Labeled 0.26
Self+RM 0.01
Table 6: Attack Success Rate comparison for IPO across different preference data creation methods.
Consistent with DPO results, Self+RM achieves the lowest ASR, while preference data from stronger models or humans leads to significantly higher ASRs. This confirms that our core conclusion extends beyond the DPO itself but also some other direct preference optimization methods like IPO.
Furthermore, methods like SimPO introduce adjustments such as length-normalized reward scoring to address length bias, as discussed in the Discussion section of our paper. While such methods may help mitigate specific artifacts (e.g., response length), they do not fully resolve the vulnerabilities tied to multi-model preference data. As demonstrated in our main results (Table 2 and Figure 3), length control alone is insufficient to improve safety when using externally sourced preferences, indicating that the fundamental issue lies in the source and quality of preference data rather than optimization-specific artifacts.
A.6 Additional Experiments for Online DPO
Our further experiments confirm that the benefits of self-generated preference data extend to the online DPO setting in terms of safety. We conducted experiments using Llama-3.1-8B under an online DPO setup for three iterations. In each iteration, we generated a 10,000 preference dataset using the current model. As shown in the Table 7, we found Self+RM continuously improves safety across iterations.
Iteration KW-ASR GPT-ASR GPT-Score
0 (SFT) 0.82 0.56 3.62
1 0.43 0.16 1.89
2 0.35 0.09 1.43
3 0.33 0.05 1.38
Table 7: Safety evaluation across online DPO iterations using Self+RM preference data creation.
A.7 Additional Experiments on Toxicity
To explore the broader applicability of our findings, we conducted additional experiments focusing on toxicity and fairness other than jailbreak as shown in the main paper.
We evaluated toxicity score using TOXIGEN (Hartvigsen et al., 2022), a comprehensive dataset covering over 135k toxic and 135k benign statements across 13 minority identity groups (e.g., African Americans, women, LGBTQ+ individuals, etc.). This evaluation framework allows us to assess whether our preference data creation methods affect model behavior toward different demographic groups and their propensity to generate toxic content.
For evaluation, we scored model outputs on a 1–5 Toxic-Score scale using GPT-4o, where higher scores indicate more toxic content. Table 8 presents the toxicity evaluation results for Llama-3.1-8B across all preference data creation methods.
Method Toxic-Score
HC+Self 1.87
GPT4o+Self 2.17
Stronger+Self 2.13
Peer+RM 1.85
All+RM 2.00
Human-Labeled 1.88
Self+RM 1.74
Table 8: Toxicity evaluation results using TOXIGEN dataset. Lower Toxic-Score indicates better performance in avoiding toxic outputs.
The toxicity evaluation results align with our main conclusions: Self+RM achieves the lowest toxicity score, while multi-model methods, such as using GPT-4o responses as chosen examples, result in much higher Toxic-Scores. These results indicate that Self+RM not only strengthens jailbreak resistance, but also mitigates toxicity and improves fairness.